OpenStudy (anonymous):

Hey guys I need help with some questions from my Algebra 1 class 1.Solve for x: −5|x + 1| = 10 2.Solve for x: |x − 2| + 10 = 12 3.The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? 4.Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. |5x − 6| = −41 |7x + 13| = 27

OpenStudy (shamil98):

Okay. Let's get started.. First you want the absolute value equation by itself and then solve for it.. -5|x+1| = 10 divide both sides by -5 |x+1| = -2 and now you have two equations. x + 1 = -2 and -(x+1)=-2 or simply x+1 = 2

OpenStudy (anonymous):

Mkay I follow

OpenStudy (shamil98):

just solve it from there. x+1 = -2 , x + 1 = 2 you get x = -3 , and x = 1. Try doing the second one.

OpenStudy (anonymous):

Okay I got x=4 and x=0

OpenStudy (shamil98):

|x − 2| = 2 x = 4 , x = 0, yep correct.

OpenStudy (anonymous):

Woohoo :D

OpenStudy (anonymous):

Can u help meh with the next one?

OpenStudy (shamil98):

2|x − 6| + 14 = 38 get |x-6| by itself first.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Hold on I think I got an answer